One mole of n2o 4 at 300 kelvin is kept in a closed container under 1 ATM. it is heated to 600 Kelvin when 20% mass of n2o 4 decomposes to nO2 the resultant pressure is?

Grade:11

1 Answers

Arun
6 years ago
Dear Sagar
 
N2O4 → 2NO2

moles of unreacted N2O4 = 1 (1 - 0.2) = 0.8 

moles of NO2 = 2 * 0.2 = 0.4
- - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - (+)
total moles n2 = 0.8 + 0.4 = 1.2

P1/(T1 n1) = P2/(T2 * n2)

1/(300 * 1) = P2/(600 * 1.2)

P2 = 2.4 atm
 
Regards
Arun (AskIITians forum expert)